what is the percent decrease from 21 to 17.43
Answer by stanbon(75887)

what is the percent decrease from 21 to 17.43
-------------
Quantity decrease = 21-17.43 = 3.57
Percent decrease:: 3.57/21 = 0.17 = 17%
==========================================
Cheerd,
Stan H.